OpenStudy (anonymous):

for the equation 2x+3y=18, find the missing value of the following ordered pairs (3,?) (-6,?) (0,?) (?,0)

Mehek (mehek14):

plug 3 for x 2*3+3y=18 3*2=6 subtract 18-6 = 12 3y = 12 divide by 3 y = 4 so the missing number for the first ordered pair is 4

Mehek (mehek14):

now we do the same for the rest

Mehek (mehek14):

plug -6 for x \(2*(-6)+3y=18\) what's 2*(-6)?

Mehek (mehek14):

@uju25 ?

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

2*(-6) = -12